### v2

# Semi-definite optimization of the measured relative entropies of quantum states and channels

Abstract
The measured relative entropies of quantum states and channels find operational significance in quantum information theory as achievable error rates in hypothesis testing tasks. They are of interest in the near term, as they correspond to hybrid quantum–classical strategies with technological requirements far less challenging to implement than required by the most general strategies allowed by quantum mechanics. In this paper, we prove that these measured relative entropies can be calculated efficiently by means of semi-definite programming, by making use of variational formulas for the measured relative entropies of states and semi-definite representations of the weighted geometric mean and the operator connection of the logarithm. Not only do the semi-definite programs output the optimal values of the measured relative entropies of states and channels, but they also provide numerical characterizations of optimal strategies for achieving them, which is of significant practical interest for designing hypothesis testing protocols.
1 Introduction
1.1 Background
The relative entropy [KL51] and its generalization to Rényi relative entropy [R6́1] are important distinguishability measures in information theory, finding direct operational meaning in hypothesis testing tasks [Che52, Ste, Che56, Str62, Hoe65, HK89] while being used to construct other entropic measures like mutual information and conditional entropy [Sib69, Aug78, Csi95]. There are a number of quantum generalizations of these quantities [Ume62, Pet85, Pet86, MLDS+13, WWY14], finding operational meaning in quantum hypothesis testing tasks [HP91, ON00, Nag06, ACM+07, Hay07, ANSV08, NS09] while being used to construct other entropic measures like quantum mutual information and conditional entropy (see, e.g., [KW20, Definition 7.18]). In these quantum hypothesis testing tasks, one assumes that there are many copies of the states available and furthermore that it is possible to perform a collective measurement on them. The technological capabilities required to perform such a collective measurement appear to be quite challenging, and it seems one might generally need a quantum computer to do so [DSM+22, CDCS24].
The measured relative entropy [Don86, Pia09] is a distinguishability measure that relaxes the requirements of quantum hypothesis testing signficantly, and it has been generalized to the Rényi family as well [Fuc96, Eqs. (3.116)–(3.117)]. Indeed, the idea behind these measures is to evaluate classical distinguishability measures on the distributions that result from performing a measurement on a single copy of the state and then optimize them over all possible measurements. See Definition 1 and Definition 2 for precise definitions of the standard and Rényi measured relative entropies, respectively. In such a way, these quantities lead to technologically feasible strategies for quantum hypothesis testing, which consist of a hybrid approach involving quantum measurement and classical post-processing. Indeed, even though there are gaps between the fundamental error rates of quantum hypothesis testing under general, collective measurements and those that result from these hybrid quantum–classical strategies, the latter strategies are more feasible in the near term.
Beyond distinguishing states, one can also distinguish quantum channels from one another, a task known as quantum channel discrimination, which has been studied extensively in quantum information [CDP08, DFY09, PW09, HHLW10, MPW10, CMW16, WW19, WBHK20, FFRS20, FF21a, BDSW24, SHW22, BKSD23]. The most general strategy allowed by quantum mechanics in such a scenario is rather complex (see [WBHK20, Figure 1]), and the technological requirements for realizing such a general strategy are even more challenging than those needed to perform a collective measurement (i.e., one would need more complex quantum computations to realize such strategies). As such, one can also consider relaxing the technological requirements for channel discrimination by considering measured relative entropies of channels, as a special case of the generalized channel divergences defined in [LKDW18, Definition II.2]. See Definition 3 and Definition 4 for precise definitions of the standard and Rényi measured relative entropies of channels, respectively, which also include energy constraints on the channel input state. Although the diamond distance [Kit97] and its energy-constrained counterpart [Shi18, Win17] are in widespread use as measures of channel distinguishability (see, e.g., [KW20, Section 6.3]), the related general notion of measured relative entropy of channels has only been explicitly defined more recently [LHT22, Eq. (8)], therein related to an operational task called sequential channel discrimination. Here we also explicitly define the measured Rényi relative entropy of channels, as a special case of the general concept from [LKDW18, Definition II.2] and [SWAT18, Eq. (12.12)].
1.2 Summary of results
In this paper, we prove that the measured relative entropies of quantum states and channels can be computed by means of semi-definite optimization algorithms (also known as semi-definite programs). These algorithms have runtimes that scale efficiently with the dimension of the states and the input and output dimensions of the channels, by employing known techniques for solving semi-definite programs [PW00, AHK05, AHK12, LSW15]. Furthermore, an added benefit of these algorithms is that, not only does one obtain the optimal values of the measured relative entropies, but one also obtains numerically an optimal measurement for the measured relative entropies of states and an optimal input state and measurement for the measured relative entropies of channels. This latter capability is of significant value for applications, in which one wishes to construct a hybrid quantum-classical strategy for achieving the error rates of hypothesis testing achievable by the measured relative entropies.
Our claims build upon two papers, which, coincidentally, were initially released on the quant-ph arXiv within two days of each other [BFT17, FS17]. Another edifice for our claims is [FSP19]. In more detail, the paper [BFT17] established variational formulas for the measured relative entropy and measured Rényi relative entropy, while the paper [FS17] proved that the hypograph and epigraph of the weighted geometric mean have efficient semi-definite representations (here, see also [Sag13]), and the paper [FSP19] proved that the hypograph of the operator connection of the logarithm has an efficient semi-definite representation. Here, we essentially combine these findings to arrive at our claims.
Indeed, for quantum states, our main contributions are to establish reductions of the variational formulas of [BFT17] to semi-definite optimization problems involving linear objective functions and the aforementioned hypographs or epigraphs (see Propositions 3 and 7). This finding is admittedly a rather direct combination of the contributions of [BFT17, FS17, FSP19]. However, it is ultimately useful in establishing our next contribution, which is an extension of these findings to measured relative entropies of channels. To establish these latter results, we use basic properties of weighted geometric means and the operator connection of the logarithm (see Propositions 10 and 12).
One benefit of our findings is that they lead to semi-definite programs involving linear matrix inequalities each of size $2d\times 2d$ when the states are $d\times d$ matrices and of size $2d_{A}d_{B}\times 2d_{A}d_{B}$ when the channels have input dimension $d_{A}$ and output dimension $d_{B}$ (see Propositions 3, 7, 10, and 12 for precise statements). As such, they do not suffer from the quadratic increase in size that occurs when applying the approach from [FS17, FSP19] to the Petz–Rényi and standard quantum relative entropies (however, note that there has been progress on addressing this issue more recently [FS23, FF23, KS24]). Furthermore, it is unclear how to apply the approach from [FS17, FSP19] for computing the dynamical (channel) version of these quantities. However, one of our main contributions is semi-definite programs for the measured relative entropies of channels, and the transition from our claims for states to our claims for channels is smooth, with the proofs consisting of just a few lines (see (71)–(80) and (89)–(98) for these steps).
1.3 Organization of the paper
The rest of our paper is organized as follows. Section 2 establishes notation and reviews background material, including the weighted geometric mean and its properties, its hypograph and epigraph, and operator connections and their properties (especially for the logarithm). The remaining Sections 3, 4, 5, and 6 provide essential definitions and detail our main results for measured Rényi relative entropy of states, measured relative entropy of states, measured Rényi relative entropy of channels, and measured relative entropy of channels, respectively. We conclude in Section 7 with a brief summary and some directions for future research.
2 Notation and Preliminaries
For a Hilbert space $\mathcal{H}$, we employ the following notation:
$\mathbb{L}(\mathcal{H})$ set of linear operators acting on $\mathcal{H}$
$\mathbb{H}(\mathcal{H})$ set of Hermitian operators acting on $\mathcal{H}$
$\mathbb{P}(\mathcal{H)}$ set of positive semi-definite operators acting on $\mathcal{H}$
$\mathbb{P}_{>0}(\mathcal{H)}$ set of positive definite operators acting on $\mathcal{H}$
$\mathbb{D}(\mathcal{H)}$ set of density operators acting on $\mathcal{H}$
Note that $\mathbb{D}(\mathcal{H})\coloneqq\left\{\rho\in\mathbb{P}(\mathcal{H}):\operatorname{Tr}[\rho]=1\right\}$.
A quantum channel is a completely positive and trace-preserving map that takes $\mathbb{L}(\mathcal{H})$ to $\mathbb{L}(\mathcal{K})$, where $\mathcal{K}$ is another Hilbert space. We often denote a quantum channel by $\mathcal{N}_{A\rightarrow B}$, which indicates that the input space is $\mathbb{L}(\mathcal{H}_{A})$ and the output space is $\mathbb{L}(\mathcal{H}_{B})$. See [Wil17, KW20] for further background on quantum information theory.
2.1 Weighted geometric mean and its properties
Given positive definite operators $X,Y\in\mathbb{P}_{>0}(\mathcal{H)}$, the weighted (operator) geometric mean $X\#_{t}Y$ of weight $t\in\mathbb{R}$ is defined as [PW75, KA80]
$X\#_{t}Y\coloneqq X^{1/2}\left(X^{-1/2}YX^{-1/2}\right)^{t}X^{1/2}.$ (1)
It is alternatively denoted by
$G_{t}(X,Y)\coloneqq X\#_{t}Y,$ (2)
and we adopt this notation in what follows. The following identity holds for all $t\in\mathbb{R}$ (see, e.g., [KW20, Eq. (7.6.5)]):
$G_{t}(X,Y)=G_{1-t}(Y,X),$ (3)
and so does the following identity for all $s,t\in\mathbb{R}$:
$G_{s}(X,G_{t}(X,Y))=G_{st}(X,Y).$ (4)
The function $x\mapsto x^{t}$ is operator concave and operator monotone for $t\in\left[0,1\right]$, operator antimonotone and operator convex for $t\in\left[-1,0\right]$, and operator convex for $t\in\left[1,2\right]$ (see, e.g., [Car10, Theorem 2.6]). The function
$\left(X,Y\right)\mapsto G_{t}(X,Y)$ (5)
is operator concave for $t\in\left[0,1\right]$ and operator convex for $t\in\left[-1,0\right]\cup\left[1,2\right]$. For $t\in\left[-1,1\right]$, this statement is a consequence of [KA80, Theorem 3.5] and, for $t\in\left[1,2\right]$, it is a consequence of (3) and [KA80, Theorem 3.5], as well as the aforementioned operator monotonicity properties of $x\mapsto x^{t}$. Concavity and convexity of the function $\left(X,Y\right)\mapsto G_{t}(X,Y)$ is also known as joint concavity and joint convexity of the weighted geometric mean.
A useful property of the weighted geometric mean for $t\in\left[-1,2\right]$ is the transformer inequality [KA80, Theorem 3.5]. For a linear operator $K\in\mathbb{L}(\mathcal{H})$, the following inequality holds for all $t\in\left[0,1\right]$:
$KG_{t}(X,Y)K^{{\dagger}}\leq G_{t}(KXK^{{\dagger}},KYK^{{\dagger}}),$ (6)
and the opposite inequality holds for all $t\in\left[-1,0\right]\cup\left[1,2\right]$:
$KG_{t}(X,Y)K^{{\dagger}}\geq G_{t}(KXK^{{\dagger}},KYK^{{\dagger}}).$ (7)
These inequalities are saturated when $K$ is invertible; i.e., for all $t\in\left[-1,2\right]$ and invertible $K$, the following holds:
$KG_{t}(X,Y)K^{{\dagger}}=G_{t}(KXK^{{\dagger}},KYK^{{\dagger}}).$ (8)
The inequalities in (6)–(7) were proven for all $t\in\left[-1,1\right]$ in [KA80, Theorem 3.5], and the extension to $t\in\left[1,2\right]$ follows from (3) and [KA80, Theorem 3.5]. See also [FF21a, Lemma 47].
2.2 Hypograph and epigraph of the weighted geometric mean
For $t\in\left[0,1\right]$, the operator hypograph of $G_{t}$ is given by [FS17, Section 3.1]
$\text{hyp}_{t}\coloneqq\left\{\left(X,Y,T\right)\in\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{H}(\mathcal{H}):G_{t}(X,Y)\geq T\right\},$ (9)
and for $t\in\left[-1,0\right]\cup\left[1,2\right]$, the operator epigraph of $G_{t}$ is given by [FS17, Section 3.1]
$\text{epi}_{t}\coloneqq\left\{\left(X,Y,T\right)\in\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{H}(\mathcal{H}):G_{t}(X,Y)\leq T\right\}.$ (10)
These sets are convex due to the aforementioned concavity and convexity properties of $G_{t}$.
As a consequence of [FS17, Theorem 3], for all rational $t\in\left[0,1\right]$, the set hypt is semi-definite representable (see also [Sag13]), and for all rational $t\in\left[-1,0\right]\cup\left[1,2\right]$, the set epit is semi-definite representable. This means that these sets can be represented in terms of a finite number of linear matrix inequalities [Nie12] and implies that one can use the methods of semi-definite programming to optimize over elements of these sets. This fact was put to use in [FF21a, FF21b] for quantum information-theoretic applications, and we make use of it here as well.
2.3 Operator connections
Generalizing the notion of an operator geometric mean, an operator connection is defined in terms of an operator monotone function $f$ as [KA80]
$P_{f}(X,Y)\coloneqq X^{1/2}f\!\left(X^{-1/2}YX^{-1/2}\right)X^{1/2},$ (11)
where $X,Y\in\mathbb{P}_{>0}(\mathcal{H)}$. This is also known as a non-commutative perspective function [Eff09, ENG11, EH14]. Due to [KA80, Theorem 3.5], the function
$\left(X,Y\right)\mapsto P_{f}(X,Y)$ (12)
is operator concave (i.e., jointly concave), and the transformer inequality holds for every linear operator $K\in\mathbb{L}(\mathcal{H})$:
$KP_{f}(X,Y)K^{{\dagger}}\leq P_{f}(KXK^{{\dagger}},KYK^{{\dagger}}).$ (13)
Equality holds in (13) if $K$ is invertible; i.e., for invertible $K\in\mathbb{L}(\mathcal{H})$, the following equality holds:
$KP_{f}(X,Y)K^{{\dagger}}=P_{f}(KXK^{{\dagger}},KYK^{{\dagger}}).$ (14)
The operator hypograph of $P_{f}$ is given by
$\text{hyp}_{f}\coloneqq\left\{\left(X,Y,T\right)\in\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{P}_{>0}(\mathcal{H)\times}\mathbb{H}(\mathcal{H}):P_{f}(X,Y)\geq T\right\},$ (15)
and it is a convex set due to the aforementioned joint concavity of $P_{f}(X,Y)$.
The logarithm is the main example of an operator monotone function on which we focus, other than the power functions from Section 2.1 , due to its connection with relative entropy. Furthermore, there is an efficient semi-definite approximation of the hypograph of the connection of the logarithm (i.e., hypln[FSP19, Theorem 3], which leads to semi-definite optimization algorithms for calculating measured relative entropies of states and channels. This fact was put to use in [FF18, Wil18, BGKE20, CGW21, WWW24] for quantum information-theoretic applications, and we make use of it here as well. To be clear, we use the following notation later on:
$P_{\ln}(X,Y)\coloneqq X^{1/2}\ln\!\left(X^{-1/2}YX^{-1/2}\right)X^{1/2}.$ (16)
This is also related to the operator relative entropy [FK89] , for which one finds the following notation in the literature [FF21a]:
$D_{\text{op}}(X\|Y)$ $\coloneqq-P_{\ln}(X,Y)$ (17)
$=X^{1/2}\ln\!\left(X^{1/2}Y^{-1}X^{1/2}\right)X^{1/2}.$ (18)
3 Measured Rényi relative entropy of states
3.1 Definition and basic properties
Given a probability distribution $p\equiv\left(p(x)\right)_{x\in\mathcal{X}}$ and a non-negative function $q\equiv\left(q(x)\right)_{x\in\mathcal{X}}$, the Rényi relative entropy is defined for $\alpha\in\left(0,1\right)\cup(1,\infty)$ as [R6́1]
$D_{\alpha}(p\|q)\coloneqq\frac{1}{\alpha-1}\ln\sum_{x\in\mathcal{X}}p(x)^{\alpha}q(x)^{1-\alpha},$ (19)
when $\alpha\in\left(0,1\right)$ or when $\alpha>1$ and $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$. Otherwise, when $\alpha>1$ and $\operatorname{supp}(p)\not\subseteq\operatorname{supp}(q)$, it is set to $\infty$. The Rényi relative entropy satisfies the data-processing inequality for all $\alpha\in\left(0,1\right)\cup(1,\infty)$, which means that
$D_{\alpha}(p\|q)\geq D_{\alpha}(N(p)\|N(q)),$ (20)
where $N$ is a classical channel (i.e., a conditional probability distribution with elements $\left(N(y|x)\right)_{y\in\mathcal{Y},x\in\mathcal{X}}$) and the notation $N(p)$ is a shorthand for the distribution that results from processing $p$ with $N$:
$N(p)\equiv\left(\sum_{x\in\mathcal{X}}N(y|x)p(x)\right)_{y\in\mathcal{Y}},$ (21)
with a similar meaning for $N(q)$. The Rényi relative entropy is also monotone in the parameter $\alpha$; i.e., for $\beta>\alpha>0$, the following inequality holds for all $p$ and $q$:
$D_{\alpha}(p\|q)\leq D_{\beta}(p\|q).$ (22)
See [vEH14] for a review of the classical Rényi relative entropy in (19).
Definition 1 (Measured Rényi relative entropy)
Given a quantum state $\rho$ and a positive semi-definite operator $\sigma$, the measured Rényi relative entropy is defined by optimizing the Rényi relative entropy over all possible measurements [Fuc96, Eqs. (3.116)–(3.117)]:
$D_{\alpha}^{M}(\rho\|\sigma)\coloneqq\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\frac{1}{\alpha-1}\ln\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Lambda_{x}\rho]^{\alpha}\operatorname{Tr}[\Lambda_{x}\sigma]^{1-\alpha},$ (23)
where the supremum is over every finite alphabet $\mathcal{X}$ and every positive operator-valued measure (POVM) $\left(\Lambda_{x}\right)_{x\in\mathcal{X}}$ (i.e., satisfying $\Lambda_{x}\geq 0$ for all $x\in\mathcal{X}$ and $\sum_{x\in\mathcal{X}}\Lambda_{x}=I$).
For $\alpha>1$, the measured Rényi relative entropy is finite if and only if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$. If the support condition holds, then it follows that the support of $\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}$ is contained in the support of $\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}}$, which in turn implies that $D_{\alpha}^{M}(\rho\|\sigma)<+\infty$. If the support condition does not hold, then $D_{\alpha}^{M}(\rho\|\sigma)=+\infty$, by applying the argument in [RSB24, Section 3.5].
We now recall some basic properties of the measured Rényi relative entropies, the first of which is actually a consequence of [BFT17, Theorem 4] and the second observed in [RSB24, Lemma 5].
Proposition 1
It suffices to optimize $D_{\alpha}^{M}(\rho\|\sigma)$ over rank-one POVMs; i.e.,
$D_{\alpha}^{M}(\rho\|\sigma)=\sup_{\mathcal{X},\left(\varphi_{x}\right)_{x\in\mathcal{X}}}\frac{1}{\alpha-1}\ln\sum_{x\in\mathcal{X}}\operatorname{Tr}[\varphi_{x}\rho]^{\alpha}\operatorname{Tr}[\varphi_{x}\sigma]^{1-\alpha},$ (24)
where each $\varphi_{x}$ is a rank-one operator such that $\sum_{x\in\mathcal{X}}\varphi_{x}=I$.
Proof. This is a direct consequence of the data-processing inequality in (20). Indeed, by diagonalizing $\Lambda_{x}$ as $\Lambda_{x}=\sum_{z\in\mathcal{Z}}\phi_{x,z}$, where each $\phi_{x,z}$ is rank one, consider that every POVM $\left(\Lambda_{x}\right)_{x\in\mathcal{X}}$ can be understood as a coarse graining of the POVM $\left(\phi_{x,z}\right)_{x\in\mathcal{X},z\in\mathcal{Z}}$ because
$\operatorname{Tr}[\Lambda_{x}\rho]=\sum_{z\in\mathcal{Z}}\operatorname{Tr}[\phi_{x,z}\rho].$ (25)
By defining $p_{X,Z}(x,z)\coloneqq\operatorname{Tr}[\phi_{x,z}\rho]$ and $q_{X,Z}(x,z)\coloneqq\operatorname{Tr}[\phi_{x,z}\sigma]$ and noting that one obtains $p_{X}(x)=\operatorname{Tr}[\Lambda_{x}\rho]$ and $q_{X}(x)=\operatorname{Tr}[\Lambda_{x}\sigma]$ by marginalization (a particular kind of classical channel), the data-processing inequality in (20) implies that
$D_{\alpha}(p_{X,Z}\|q_{X,Z})\geq D_{\alpha}(p_{X}\|q_{X}),$ (26)
concluding the proof.
Proposition 2
The measured Rényi relative entropy obeys the data-processing inequality; i.e., for every state $\rho$, positive semi-definite operator $\sigma$, quantum channel $\mathcal{N}$, and $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$, the following inequality holds:
$D_{\alpha}^{M}(\rho\|\sigma)\geq D_{\alpha}^{M}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)).$ (27)
Proof. Observe that
$\frac{1}{\alpha-1}\ln\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Lambda_{x}\mathcal{N}(\rho)]^{\alpha}\operatorname{Tr}[\Lambda_{x}\mathcal{N}(\sigma)]^{1-\alpha}$
$=\frac{1}{\alpha-1}\ln\sum_{x\in\mathcal{X}}\operatorname{Tr}[\mathcal{N}^{{\dagger}}(\Lambda_{x})\rho]^{\alpha}\operatorname{Tr}[\mathcal{N}^{{\dagger}}(\Lambda_{x})\sigma]^{1-\alpha}$ (28)
$\leq D_{\alpha}^{M}(\rho\|\sigma).$ (29)
In the above, we made use of the Hilbert–Schmidt adjoint $\mathcal{N}^{{\dagger}}$, which is completely positive and unital, implying that $\left(\mathcal{N}^{{\dagger}}(\Lambda_{x})\right)_{x\in\mathcal{X}}$ is a POVM. The inequality follows from the fact that $D_{\alpha}^{M}(\rho\|\sigma)$ involves an optimization over every alphabet $\mathcal{X}$ and POVM. Since the inequality holds for every POVM $\left(\Lambda_{x}\right)_{x\in\mathcal{X}}$, we conclude (27). (Here we can also observe that the claim holds more generally for positive, trace-preserving maps.)
It can be helpful to write the measured Rényi relative entropy in terms of the measured Rényi relative quasi-entropy:
$D_{\alpha}^{M}(\rho\|\sigma)=\frac{1}{\alpha-1}\ln Q_{\alpha}^{M}(\rho\|\sigma),$ (30)
where the latter is defined as
$Q_{\alpha}^{M}(\rho\|\sigma)\coloneqq\left\{\begin{array}[c]{cc}\inf_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Lambda_{x}\rho]^{\alpha}\operatorname{Tr}[\Lambda_{x}\sigma]^{1-\alpha}&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Lambda_{x}\rho]^{\alpha}\operatorname{Tr}[\Lambda_{x}\sigma]^{1-\alpha}&\text{for }\alpha>1\end{array}\right.$. (31)
One can also define the projectively measured Rényi relative entropy as
$D_{\alpha}^{P}(\rho\|\sigma)\coloneqq\frac{1}{\alpha-1}\ln Q_{\alpha}^{P}(\rho\|\sigma),$ (32)
where
$Q_{\alpha}^{P}(\rho\|\sigma)\coloneqq\left\{\begin{array}[c]{cc}\inf_{\left(\Pi_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Pi_{x}\rho]^{\alpha}\operatorname{Tr}[\Pi_{x}\sigma]^{1-\alpha}&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\left(\Pi_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Pi_{x}\rho]^{\alpha}\operatorname{Tr}[\Pi_{x}\sigma]^{1-\alpha}&\text{for }\alpha>1\end{array}\right$., (33)
with the key difference being that the optimization is performed over every projective measurement $\left(\Pi_{x}\right)_{x\in\mathcal{X}}$ (i.e., satisfying $\Pi_{x}\Pi_{x^{\prime}}=\Pi_{x}\delta_{x,x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{X}$ in addition to the requirements of a POVM) and the size of the alphabet $\mathcal{X}$ is equal to the dimension of the underlying Hilbert space of $\rho$ and $\sigma$.
It is known from [BFT17, Theorem 4] that the following equalities hold for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$:
$D_{\alpha}^{M}(\rho\|\sigma)=D_{\alpha}^{P}(\rho\|\sigma),\qquad Q_{\alpha}^{M}(\rho\|\sigma)=Q_{\alpha}^{P}(\rho\|\sigma),$ (34)
which is a non-trivial finding that makes use of operator concavity and convexity properties of the function $x\mapsto x^{t}$. Furthemore, it was noted therein that the measured Rényi relative entropy is achieved by a rank-one, projective measurement. This has practical implications for achieving the measured Rényi relative entropy because projective measurements are simpler to realize experimentally than general POVMs.
3.2 Variational formulas for measured Rényi relative entropy of states
As a consequence of [BFT17, Lemma 3 and Theorem 4], the measured Rényi relative entropy has the following variational formulas for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$:
$D_{\alpha}^{M}(\rho\|\sigma)$ $=\sup_{\omega>0}\left\{\frac{1}{\alpha-1}\ln\!\left(\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right)\right\}$ (35)
$=\sup_{\omega>0}\left\{\frac{1}{\alpha-1}\ln\!\left(\left(\operatorname{Tr}[\omega\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right)^{1-\alpha}\right)\right\}.$ (36)
These are a direct consequence of and equivalent to the precise expressions given in [BFT17, Lemma 3], which are as follows:
$Q_{\alpha}^{M}(\rho\|\sigma)$ $=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right\}&\text{for }\alpha\in\left(0,1/2\right)\\ \inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha\in[1/2,1)\\ \sup_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha>1\end{array}\right$., (40)
$Q_{\alpha}^{M}(\rho\|\sigma)$ $=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\left(\operatorname{Tr}[\omega\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\omega>0}\left\{\left(\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha>1\end{array}\right.$. (43)
Indeed, one obtains (35) for $\alpha\geq 1/2$ from the second and third expressions in (40) by the substitution $\omega\rightarrow\omega^{\frac{\alpha}{\alpha-1}}$, and similarly for getting (36) from (43) for $\alpha>1$. We note in passing that these variational formulas have found application in devising variational quantum algorithms for estimating the measured Rényi relative entropy [GPSW24].
Although the expressions in (35)–(36) are simpler than those in (40)–(43), the various expressions in (40) are helpful for seeing that the optimizations can be performed efficiently (see Proposition 3 for further details). To see this, let us consider the expressions in (40) one at a time. For $\alpha\in\left(0,1/2\right)$, the function $\omega\mapsto\omega^{\frac{\alpha}{\alpha-1}}$ is operator convex because $\frac{\alpha}{\alpha-1}\in\left(-1,0\right)$. As such, the objective function $\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]$ is convex in $\omega$. For $\alpha\in[1/2,1)$, the function $\omega\mapsto\omega^{1-\frac{1}{\alpha}}$ is operator convex because $1-\frac{1}{\alpha}\in[-1,0)$. Then the objective function $\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]$ is convex in $\omega$. Finally, for $\alpha>1$, the function $\omega\mapsto\omega^{1-\frac{1}{\alpha}}$ is operator concave because $1-\frac{1}{\alpha}\in\left(0,1\right)$. Then the objective function $\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]$ is concave in $\omega$.
An operator $\omega$ that achieves the optimal values in (40)–(43) corresponds to an observable whose eigenvectors form an optimal measurement for achieving the measured Rényi relative entropy. This point becomes clear by inspecting the proof of Proposition 13 below. As such, being able to calculate such an observable numerically is valuable from an operational perspective, and we note here that this task is accomplished by the semi-definite optimization algorithm mentioned in Proposition 3. It has been known for some time that the optimal observable for $\alpha=1/2$ has an analytical form [FC95], given by $G_{\frac{1}{2}}(\sigma^{-1},\rho)$ and known as the Fuchs–Caves observable (see also [Fuc96, Section 3.3]).
In Appendix A, we provide an alternative proof of (35), which makes use of the inequality of arithmetic and geometric means, as well as Bernoulli’s inequality. We think this proof is of interest due to its simplicity. Let us note that the expressions in (43) were actually shown in the proof of [BFT17, Lemma 3] to follow from (40) by means of these inequalities.
3.3 Optimizing the measured Rényi relative entropy of states
One of the main goals of [BFT17] was to derive variational formulas for the measured Rényi relative entropy and explore applications of them in quantum information. In this section, we observe in Proposition 3 below that there is an alternative variational representation of the measured Rényi relative entropy in terms of a linear objective function and the hypograph or epigraph of the weighted geometric mean. From this observation, we conclude that there is an efficient semi-definite optimization algorithm for computing the measured Rényi relative entropy, which makes use of the fact recalled in Section 2.2 (i.e., from [FS17, Theorem 3]). As mentioned previously, not only does this algorithm compute the optimal value of $Q_{\alpha}^{M}(\rho\|\sigma)$ for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$, but it also determines an optimal observable $\omega$. Another advantage of the variational representations in Proposition 3 is that they lead to a rather rapid derivation of variational representations of the measured Rényi relative entropy of channels (see Proposition 10).
Proposition 3
Let $\rho$ be a state and $\sigma$ a positive semi-definite operator. For $\alpha\in\left(0,1/2\right)$,
$Q_{\alpha}^{M}(\rho\|\sigma)=\inf_{\omega,\theta>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\theta\sigma]:\theta\geq G_{\frac{\alpha}{\alpha-1}}(I,\omega)\right\},$ (44)
for $\alpha\in[1/2,1)$,
$Q_{\alpha}^{M}(\rho\|\sigma)=\inf_{\omega,\theta>0}\left\{\alpha\operatorname{Tr}[\theta\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]:\theta\geq G_{1-\frac{1}{\alpha}}(I,\omega)\right\},$ (45)
and for $\alpha>1$,
$Q_{\alpha}^{M}(\rho\|\sigma)=\sup_{\omega,\theta>0}\left\{\alpha\operatorname{Tr}[\theta\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]:\theta\leq G_{1-\frac{1}{\alpha}}(I,\omega)\right\}.$ (46)
For all rational $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$, the quantity $Q_{\alpha}^{M}(\rho\|\sigma)$ can be calculated by means of a semi-definite program. More specifically, when $\rho$ and $\sigma$ are $d\times d$ matrices and $p$ and $q$ are relatively prime integers such that $\frac{p}{q}=\frac{\alpha}{\alpha-1}$ for $\alpha\in\left(0,1/2\right)$ or $\frac{p}{q}=1-\frac{1}{\alpha}$ for $\alpha\in\left[1/2,1\right)\cup(1,\infty)$, the semi-definite program requires $O(\log_{2}q)$ linear matrix inequalities each of size $2d\times 2d$.
Proof. These formulas are a direct consequence of (40) and the following identities:
$\omega^{\frac{\alpha}{\alpha-1}}=G_{\frac{\alpha}{\alpha-1}}(I,\omega),\qquad\omega^{1-\frac{1}{\alpha}}=G_{1-\frac{1}{\alpha}}(I,\omega),$ (47)
while noting that the optimal value of $\theta$ in (44) is equal to $G_{\frac{\alpha}{\alpha-1}}(I,\omega)$ and the optimal value of $\theta$ in (45) and (46) is equal to $G_{1-\frac{1}{\alpha}}(I,\omega)$.
As such, we have rewritten $Q_{\alpha}^{M}(\rho\|\sigma)$ for $\alpha\in\left(0,1/2\right)$ in terms of the hypograph of $G_{\frac{\alpha}{\alpha-1}}$, for $\alpha\in[1/2,1)$ in terms of the hypograph of $G_{1-\frac{1}{\alpha}}$, and for $\alpha>1$ in terms of the epigraph of $G_{1-\frac{1}{\alpha}}$. By appealing to [FS17, Theorem 3], it follows that all of these quantities can be efficiently calculated for rational $\alpha$ by means of semi-definite programming, with complexity as stated above.
4 Measured relative entropy of states
4.1 Definition and basic properties
Given a probability distribution $p\equiv\left(p(x)\right)_{x\in\mathcal{X}}$ and a non-negative function $q\equiv\left(q(x)\right)_{x\in\mathcal{X}}$, the relative entropy is defined as
$D(p\|q)\coloneqq\sum_{x\in\mathcal{X}}p(x)\ln\!\left(\frac{p(x)}{q(x)}\right),$ (48)
when $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$ and it is set to $\infty$ otherwise. It is equal to the $\alpha\rightarrow 1$ limit of the Rényi relative entropy in (19):
$D(p\|q)=\lim_{\alpha\rightarrow 1}D_{\alpha}(p\|q).$ (49)
By virtue of the ordering property in (22), we can write
$D(p\|q)=\sup_{\alpha\in\left(0,1\right)}D_{\alpha}(p\|q)=\inf_{\alpha>1}D_{\alpha}(p\|q).$ (50)
As a direct consequence of (20) and (49), the relative entropy satisfies the data-processing inequality, which means that
$D(p\|q)\geq D(N(p)\|N(q)),$ (51)
where we used the same notation from (21).
Definition 2 (Measured relative entropy)
Given a quantum state $\rho$ and a positive semi-definite operator $\sigma$, the measured relative entropy is defined by optimizing the relative entropy over all possible measurements [Don86, Pia09]:
$D^{M}(\rho\|\sigma)\coloneqq\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\Lambda_{x}\rho]\ln\!\left(\frac{\operatorname{Tr}[\Lambda_{x}\rho]}{\operatorname{Tr}[\Lambda_{x}\sigma]}\right),$ (52)
where the supremum is over every finite alphabet $\mathcal{X}$ and every positive operator-valued measure (POVM) $\left(\Lambda_{x}\right)_{x\in\mathcal{X}}$ (i.e., satisfying $\Lambda_{x}\geq 0$ for all $x\in\mathcal{X}$ and $\sum_{x\in\mathcal{X}}\Lambda_{x}=I$).
Proposition 4
For $\rho$ a state and $\sigma$ a positive semi-definite operator, the measured relative entropy is equal to the $\alpha\rightarrow 1$ limit of the measured Rényi relative entropy:
$D^{M}(\rho\|\sigma)=\lim_{\alpha\rightarrow 1}D_{\alpha}^{M}(\rho\|\sigma).$ (53)
Proof. See Appendix B.
Let us note that the convergence statement above can be made more precise by invoking (22) and [CDL+25, Lemma 23] (see also [BDSW24, Lemma 13]). Specifically, when $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, there exists a state-dependent constant $c(\rho\|\sigma)$ such that, for all $\delta\in(0,\frac{\ln 3}{2c(\rho\|\sigma)}]$, the following bound holds:
$D_{1-\delta}^{M}(\rho\|\sigma)\leq D^{M}(\rho\|\sigma)\leq D_{1-\delta}^{M}(\rho\|\sigma)+\delta K\left[c(\rho\|\sigma)\right]^{2},$ (54)
where $K\coloneqq\cosh(\left(\ln 3\right)/2)$. We invoke this bound later on in Remark 8.
The following statements can be proven similarly to Propositions 1 and 2, or alternatively, they can be understood as the $\alpha\rightarrow 1$ limit of these propositions.
Proposition 5
It suffices to optimize $D^{M}(\rho\|\sigma)$ over rank-one POVMs; i.e.,
$D^{M}(\rho\|\sigma)=\sup_{\mathcal{X},\left(\varphi_{x}\right)_{x\in\mathcal{X}}}\sum_{x\in\mathcal{X}}\operatorname{Tr}[\varphi_{x}\rho]\ln\!\left(\frac{\operatorname{Tr}[\varphi_{x}\rho]}{\operatorname{Tr}[\varphi_{x}\sigma]}\right),$ (55)
where each $\varphi_{x}$ is a rank-one operator such that $\sum_{x\in\mathcal{X}}\varphi_{x}=I$.
Proposition 6
The measured relative entropy obeys the data-processing inequality; i.e., for every state $\rho$, positive semi-definite operator $\sigma$, and quantum channel $\mathcal{N}$, the following inequality holds:
$D^{M}(\rho\|\sigma)\geq D^{M}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)).$ (56)
4.2 Variational formulas for the measured relative entropy of states
For a state $\rho$ and a positive semi-definite operator $\sigma$, the following variational formulas for the measured relative entropy are known:
$D^{M}(\rho\|\sigma)$ $=\sup_{\omega>0}\left\{\operatorname{Tr}[\left(\ln\omega\right)\rho]-\ln\operatorname{Tr}[\omega\sigma]\right\}$ (57)
$=\sup_{\omega>0}\left\{\operatorname{Tr}[\left(\ln\omega\right)\rho]-\operatorname{Tr}[\omega\sigma]+1\right\}.$ (58)
The first was established in [Pet07, Eq. (11.69)] and [BFT17, Lemma 1 and Theorem 2], while the second was established in [BFT17, Lemma 1 and Theorem 2]. Due to the fact that the function $\omega\mapsto\ln\omega$ is operator concave, it follows that the function $\omega\mapsto\operatorname{Tr}[\left(\ln\omega\right)\rho]-\operatorname{Tr}[\omega\sigma]+1$ is concave, which is a notable feature of the variational representation in (58) that has further implications discussed in the next section.
4.3 Optimizing the measured relative entropy of states
In this section, we observe that there is an efficient algorithm for computing the measured relative entropy, which makes use of the observation in (60) and the fact recalled in Section 2.3.
Proposition 7
Let $\rho$ be a state and $\sigma$ a positive semi-definite operator. Then
$D^{M}(\rho\|\sigma)=\sup_{\omega>0,\theta\in\operatorname{Herm}}\left\{\operatorname{Tr}[\theta\rho]-\operatorname{Tr}[\omega\sigma]+1:\theta\leq P_{\ln}(I,\omega)\right\},$ (59)
where $P_{\ln}$ is defined in (16).
Furthermore, when $\rho$ and $\sigma$ are $d\times d$ matrices, the quantity $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program up to an additive error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d\times 2d$.
Proof. The formula above is a direct consequence of (58) and the following identity:
$\ln\omega=P_{\ln}(I,\omega),$ (60)
while noting that the optimal value of $\theta$ in (59) is equal to $P_{\ln}(I,\omega)$, the latter being a Hermitian operator.
As such, we have rewritten $D^{M}(\rho\|\sigma)$ in terms of the hypograph of $P_{\ln}(I,\omega)$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that $D^{M}(\rho\|\sigma)$ can be efficiently calculated by means of a semi-definite program with the stated complexity.
Remark 8
An alternative approach for computing the measured relative entropy is to set $\alpha=1-2^{-\ell}$ for $\ell\in\mathbb{N}$, similar to what was done in [FF21a, Lemma 9]. By appealing to Proposition 3 and (54), it follows that, in order to achieve an error $\varepsilon$ in computing the measured relative entropy, the semi-definite program resulting from this approach requires $O(\ln(1/\varepsilon))$ linear matrix inequalities, each of size $2d\times 2d$. When compared to the performance of the approach from Proposition 7, it is clear that this latter approach is preferred because it requires only $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities to achieve the same error.
5 Measured Rényi relative entropy of channels
5.1 Definition and basic properties
Definition 3 (Measured Rényi channel divergence)
Given a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$ (Hermitian operator acting on system $A$), and an energy constraint $E\in\mathbb{R}$, the energy-constrained measured Rényi relative entropy of channels is defined for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$ as
$D_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})\coloneqq\\ \sup_{\begin{subarray}{c}d_{R^{\prime}}\in\mathbb{N},\\ \rho_{R^{\prime}A}\in\mathbb{D}(\mathcal{H}_{R^{\prime}A}\mathcal{)}\end{subarray}}\left\{D_{\alpha}^{M}(\mathcal{N}_{A\rightarrow B}(\rho_{R^{\prime}A})\|\mathcal{M}_{A\rightarrow B}(\rho_{R^{\prime}A})):\operatorname{Tr}[H_{A}\rho_{A}]\leq E\right\}.$ (61)
In what follows, for brevity, we also refer to the quantity in (61) as the measured Rényi channel divergence.
In (61) above, the supremum is taken not only over every bipartite state $\rho_{R^{\prime}A}$ but also over the reference system $R^{\prime}$ with dimension $d_{R^{\prime}}$. We define the measured Rényi channel divergence in this general way in order to allow for all physically feasible ways of processing a channel; indeed, one prepares a state $\rho_{R^{\prime}A}$, sends it through either $\mathcal{N}_{A\rightarrow B}$ or $\mathcal{M}_{A\rightarrow B}$, and processes the systems $R^{\prime}B$ with a measurement in order to distinguish the maps $\mathcal{N}_{A\rightarrow B}$ or $\mathcal{M}_{A\rightarrow B}$. We impose an energy constraint only on the input system $A$, because this is the simplest and most minimal modification of an unconstrained channel divergence and it allows for all physically plausible, yet unconstrained, reference systems. Imposing the energy constraint in this way furthermore has the benefit of leading to an efficient algorithm for computing $D_{\alpha,H,E}^{M}$ by semi-definite programming (see Proposition 10). Let us note that imposing an energy constraint only on the input system $A$ is similar to the approach taken when defining the Shirokov–Winter energy-constrained diamond norm [Shi18, Win17] or more general energy-constrained channel divergences [SWAT18, Eq. (12.12)]. One obtains the unconstrained measured Rényi relative entropy of channels by setting $H_{A}=I_{A}$ and $E=1$, so that the “energy constraint” becomes redundant with the constraint that $\rho_{R^{\prime}A}$ is a state.
By appealing to (31), we can also write
$D_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\frac{1}{\alpha-1}\ln Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M}),$ (62)
where
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})\coloneqq\\ \left\{\begin{array}[c]{cc}\inf_{\begin{subarray}{c}d_{R^{\prime}}\in\mathbb{N},\\ \rho_{R^{\prime}A}\in\mathbb{D}(\mathcal{H}_{R^{\prime}A}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}Q_{\alpha}^{M}(\mathcal{N}_{A\rightarrow B}(\rho_{R^{\prime}A})\|\mathcal{M}_{A\rightarrow B}(\rho_{R^{\prime}A}))&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\begin{subarray}{c}d_{R^{\prime}}\in\mathbb{N},\\ \rho_{R^{\prime}A}\in\mathbb{D}(\mathcal{H}_{R^{\prime}A}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}Q_{\alpha}^{M}(\mathcal{N}_{A\rightarrow B}(\rho_{R^{\prime}A})\|\mathcal{M}_{A\rightarrow B}(\rho_{R^{\prime}A}))&\text{for }\alpha>1\end{array}\right.$. (63)
Although the optimization in (61) is defined to be over an unbounded space, it is possible to simplify the task by employing basic quantum information-theoretic reasoning. Indeed, as stated in Proposition 9 below, one can write $D_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$ and $Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$ in terms of the Choi operators of $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$, defined as
$\Gamma_{RB}^{\mathcal{N}}\coloneqq\mathcal{N}_{A\rightarrow B}(\Gamma_{RA}),\qquad\Gamma_{RB}^{\mathcal{M}}\coloneqq\mathcal{M}_{A\rightarrow B}(\Gamma_{RA}),$ (64)
where the maximally entangled operator $\Gamma_{RA}$ is defined as
$\Gamma_{RA}\coloneqq\sum_{i,j}|i\rangle\!\langle j|_{R}\otimes|i\rangle\!\langle j|_{A},$ (65)
so that the reference system $R$ is isomorphic to the channel input system $A$ (i.e., the corresponding Hilbert spaces $\mathcal{H}_{R}$ and $\mathcal{H}_{A}$ are isomorphic, denoted by $\mathcal{H}_{R}\simeq\mathcal{H}_{A}$):
Proposition 9
Given a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$, and an energy constraint $E\in\mathbb{R}$, the measured Rényi channel divergence can be written as follows for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$:
$D_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}),\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{D_{\alpha}^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})\right\},$ (66)
where $\mathcal{H}_{R}\simeq\mathcal{H}_{A}$ and $\rho_{R}=\rho_{A}$. Equivalently,
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\\ \left\{\begin{array}[c]{cc}\inf_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}Q_{\alpha}^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}Q_{\alpha}^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})&\text{for }\alpha>1\end{array}\right$. (67)
Proof. See [CMW16, Lemma 6] or [KW20, Proposition 7.82] for a detailed proof. The only difference with the optimization above and those from  [CMW16, Lemma 6] and [KW20, Proposition 7.82] is the additional energy constraint $\operatorname{Tr}[H_{A}\rho_{A}]\leq E$, which follows because $\rho_{R}^{1/2}\Gamma_{RA}\rho_{R}^{1/2}$, with $\rho_{R}=\rho_{A}$, is the canonical purification of the state $\rho_{A}$.
5.2 Optimizing the measured Rényi relative entropy of channels
In this section, we observe in Proposition 10 that there is a variational representation of the measured Rényi channel divergence in terms of a linear objective function and the hypograph or epigraph of the weighted geometric mean. From this observation, we conclude that there is an efficient semi-definite optimization algorithm for computing the measured Rényi channel divergence, which makes use of the fact recalled in Section 2.2. Not only does this algorithm compute the optimal value of $Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$ for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$, but it also determines an optimal input state and measurement that achieves $Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$, which is of significant value for applications. The proof of Proposition 10 results from a straightforward combination of Proposition 3, Proposition 9, and the transformer equality in (8).
Proposition 10
Given is a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$, and an energy constraint $E\in\mathbb{R}$. Let $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ be the Choi operators of $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$, respectively. For $\alpha\in\left(0,1/2\right)$,
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\inf_{\Omega,\Theta,\rho>0}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\Omega\Gamma^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\Theta\Gamma^{\mathcal{M}}]:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\geq G_{\frac{\alpha}{\alpha-1}}(\rho\otimes I,\Omega)\end{array}\right\},$ (68)
for $\alpha\in[1/2,1)$,
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\inf_{\Omega,\Theta,\rho>0}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\geq G_{1-\frac{1}{\alpha}}(\rho\otimes I,\Omega)\end{array}\right\},$ (69)
and for $\alpha>1$,
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\Theta,\rho>0}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq G_{1-\frac{1}{\alpha}}(\rho\otimes I,\Omega)\end{array}\right\}.$ (70)
For rational $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$, the quantity $Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$ can be calculated by means of a semi-definite program. More specifically, when $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ are $d_{A}d_{B}\times d_{A}d_{B}$ matrices and $p$ and $q$ are relatively prime integers such that $\frac{p}{q}=\frac{\alpha}{\alpha-1}$ for $\alpha\in\left(0,1/2\right)$ or $\frac{p}{q}=1-\frac{1}{\alpha}$ for $\alpha\in\left[1/2,1\right)\cup(1,\infty)$, the semi-definite program requires $O(\log_{2}q)$ linear matrix inequalities each of size $2d_{A}d_{B}\times 2d_{A}d_{B}$.
Proof. This follows by combining Proposition 3 and Proposition 9 and employing the transformer equality in (8). Let us consider the case $\alpha\in\left(0,1/2\right)$:
$Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$
$=\inf_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}Q_{\alpha}^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})$ (71)
$=\inf_{\begin{subarray}{c}\Omega^{\prime},\Theta^{\prime}>0,\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]+\left(1-\alpha\right)\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]:\\ \Theta^{\prime}\geq G_{\frac{\alpha}{\alpha-1}}(I,\Omega^{\prime})\end{array}\right\}$ (74)
$=\inf_{\begin{subarray}{c}\Omega^{\prime},\Theta^{\prime},\rho_{R}>0,\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]:\\ \Theta^{\prime}\geq G_{\frac{\alpha}{\alpha-1}}(I,\Omega^{\prime})\end{array}\right\}$ (77)
$=\inf_{\begin{subarray}{c}\Omega,\Theta,\rho_{R}>0,\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\alpha\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{M}}]:\\ \Theta\geq G_{\frac{\alpha}{\alpha-1}}(\rho\otimes I,\Omega)\end{array}\right\}.$ (80)
The first equality follows from Proposition 9. The second equality follows from Proposition 3. The third equality follows from cyclicity of trace and the fact that the function
$\rho_{R}\mapsto\alpha\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]+\left(1-\alpha\right)\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]$ (81)
is continuous in $\rho_{R}$, so that the optimization can be performed over the set of positive definite density operators (dense in the set of all density operators). The final equality follows from defining
$\Omega\coloneqq\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2},\qquad\Theta\coloneqq\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2},$ (82)
and noting that
$\Omega^{\prime},\Theta^{\prime}>0\quad\Leftrightarrow\quad\Omega,\Theta>0,$ (83)
as well as
$\quad\Theta^{\prime}\geq G_{\frac{\alpha}{\alpha-1}}(I,\Omega^{\prime})$
$\Leftrightarrow\quad\left(\rho^{1/2}\otimes I\right)\Theta^{\prime}\left(\rho^{1/2}\otimes I\right)\geq\left(\rho^{1/2}\otimes I\right)G_{\frac{\alpha}{\alpha-1}}(I,\Omega^{\prime})\left(\rho^{1/2}\otimes I\right)$ (84)
$\Leftrightarrow\quad\Theta\geq G_{\frac{\alpha}{\alpha-1}}(\rho\otimes I,\Omega),$ (85)
with the final equality following from the definitions in (82) and the transformer equality in (8). The proofs of (69) and (70) follow similarly.
As such, we have rewritten $Q_{\alpha,H,E}^{M}(\mathcal{N}\|\mathcal{M})$ for $\alpha\in\left(0,1/2\right)$ in terms of hypograph of $G_{\frac{\alpha}{\alpha-1}}$, for $\alpha\in[1/2,1)$ in terms of the hypograph of $G_{1-\frac{1}{\alpha}}$, and for $\alpha>1$ in terms of the epigraph of $G_{1-\frac{1}{\alpha}}$. By appealing to [FS17, Theorem 3], it follows that all of these quantities can be efficiently calculated for rational $\alpha$ by means of semi-definite programming, with the stated complexity.
6 Measured relative entropy of channels
6.1 Definition and basic properties
Definition 4 (Measured relative entropy of channels)
Given a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$ (Hermitian operator acting on system $A$), and an energy constraint $E\in\mathbb{R}$, the energy-constrained measured relative entropy of channels is defined as
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})\coloneqq\\ \sup_{\begin{subarray}{c}d_{R^{\prime}}\in\mathbb{N},\\ \rho_{R^{\prime}A}\in\mathbb{D}(\mathcal{H}_{R^{\prime}A}\mathcal{)}\end{subarray}}\left\{D^{M}(\mathcal{N}_{A\rightarrow B}(\rho_{R^{\prime}A})\|\mathcal{M}_{A\rightarrow B}(\rho_{R^{\prime}A})):\operatorname{Tr}[H_{A}\rho_{A}]\leq E\right\}.$ (86)
The motivation for this definition is the same as that given after (61).
Similar to what was observed in Proposition 9, although the optimization in (86) is defined to be over an unbounded space, it is possible to simplify the optimization task as follows.
Proposition 11
Given a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$, and an energy constraint $E\in\mathbb{R}$, the measured relative entropy of channels can be written as follows:
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}),\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{D^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})\right\},$ (87)
where $\mathcal{H}_{R}\simeq\mathcal{H}_{A}$ and $\rho_{R}=\rho_{A}$.
Proof. The proof is the same as that given for Proposition 9.
6.2 Optimizing the measured relative entropy of channels
In this section, we observe in Proposition 12 that there is a variational representation of the measured relative entropy of channels in terms of a linear objective function and the hypograph of the operator connection of the logarithm. From this observation, we conclude that there is an efficient semi-definite optimization algorithm for computing the measured relative entropy of channels, which makes use of the fact recalled in Section 2.3. Not only does this algorithm compute the optimal value of $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$, but it also determines an optimal input state and measurement that achieves $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$, which, as mentioned previously, is of significant value for applications. The proof of Proposition 12 results from a straightforward combination of Proposition 7, Proposition 11, and the transformer equality in (14).
Proposition 12
Given is a quantum channel $\mathcal{N}_{A\rightarrow B}$, a completely positive map $\mathcal{M}_{A\rightarrow B}$, a Hamiltonian $H_{A}$, and an energy constraint $E\in\mathbb{R}$. Let $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ be the Choi operators of $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$, respectively. Then
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\rho>0,\Theta\in\operatorname{Herm}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]+1:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\},$ (88)
where $P_{\ln}$ is defined in (16).
Furthermore, when $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ are $d_{A}d_{B}\times d_{A}d_{B}$ matrices, the quantity $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ can be efficiently calculated by means of a semi-definite program up to an error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d_{A}d_{B}\times 2d_{A}d_{B}$.
Proof. This follows similarly to the proof of Proposition 10, and we provide the proof for completeness. Indeed, here we combine Proposition 7 and Proposition 11 and employ the transformer equality in (14). Consider that
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$
$=\sup_{{}_{\begin{subarray}{c}\rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}}D^{M}(\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}\|\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2})$ (89)
$=\sup_{\begin{subarray}{c}\Omega^{\prime}>0,\Theta^{\prime}\in\operatorname{Herm},\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]-\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (92)
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\rho_{R}>0,\Theta^{\prime}\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (95)
$=\sup_{\begin{subarray}{c}\Omega,\rho_{R}>0,\Theta\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}.$ (98)
The first equality follows from Proposition 11. The second equality follows from Proposition 7. The third equality follows from cyclicity of trace and the fact that the function
$\rho_{R}\mapsto\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1$ (99)
is continuous in $\rho_{R}$, so that the optimization can be performed over the set of positive definite density operators (dense in the set of all density operators). The final equality follows from defining
$\Omega\coloneqq\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2},\qquad\Theta\coloneqq\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2},$ (100)
and noting that
$\Omega^{\prime}>0,\quad\Theta^{\prime}\in\operatorname{Herm}\quad\Leftrightarrow\quad\Omega>0,\quad\Theta\in\operatorname{Herm},$ (101)
as well as
$\quad\Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})$
$\Leftrightarrow\quad\left(\rho^{1/2}\otimes I\right)\Theta^{\prime}\left(\rho^{1/2}\otimes I\right)\leq\left(\rho^{1/2}\otimes I\right)P_{\ln}(I,\Omega^{\prime})\left(\rho^{1/2}\otimes I\right)$ (102)
$\Leftrightarrow\quad\Theta\leq P_{\ln}(\rho\otimes I,\Omega),$ (103)
with the final equality following from the definitions in (100) and the transformer equality in (14).
As such, we have rewritten $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ in terms of the hypograph of $P_{\ln}$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that this quantity can be efficiently calculated by means of semi-definite programming, with the complexity stated above.
7 Conclusion
The main contributions of our paper are efficient semi-definite optimization algorithms for computing measured relative entropies of quantum states and channels. We did so by combining the results of [BFT17, FS17, FSP19] to obtain the result for states and then we generalized these to quantum channels by using basic properties of the weighted geometric mean and operator connection of the logarithm. Our findings are of significant value for applications, in which one wishes to find numerical characterizations of technologically feasible, hybrid quantum-classical strategies for quantum hypothesis testing of states and channels.
Going forward from here, we note that further work in this direction could consider combining our findings here with those of [RSB24], the latter being about measured Rényi divergences under restricted forms of measurements. We also think it is a very interesting open question, related to those from [FS17, FS23], to determine semi-definite programs for various Rényi relative entropies of quantum channels, including those based on the sandwiched [MLDS+13, WWY14] and Petz–Rényi [Pet85, Pet86]  relative entropies. More generally, one could consider the same question for $\alpha$-$z$ Rényi relative entropies [AD15], and here we think the variational formulas from [MH24, Appendix B] could be useful in addressing this question.
Acknolwedgements—We are grateful to Ludovico Lami for communicating to us that the root fidelity of states $\rho$ and $\sigma$ can be written as $\frac{1}{2}\inf_{Y,Z>0}\left\{\operatorname{Tr}[Y\rho]+\operatorname{Tr}[Z\sigma]:G_{1/2}(Y,Z)=I\right\}$, which was a starting point for our work here. We also thank James Saunderson for several email exchanges regarding semi-definite optimization and the weighted geometric mean, related to [FS17]. ZH is supported by a Sydney Quantum Academy Postdoctoral Fellowship and an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging”. She is also grateful to the Cornell School of Electrical and Computer Engineering and the Cornell Lab of Ornithology for hospitality during a June 2024 research visit. MMW acknowledges support from the National Science Foundation under Grant No. 2304816 and from Air Force Research Laboratory under agreement number FA8750-23-2-0031.
This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.
Appendix A Alternative proof of variationals formulas for measured Rényi relative entropies
Proposition 13 ([BFT17])
For a state $\rho$ and a positive semi-definite operator $\sigma$, the expression in (35) holds.
Proof. Let us begin with the case $\alpha\in\left(0,1\right)$. Let us write an arbitrary $\omega>0$ in terms of a spectral decomposition as
$\omega=\sum_{y}\omega_{y}|\phi_{y}\rangle\!\langle\phi_{y}|,$ (104)
where $\omega_{y}>0$ for all $y$ and $\left\{|\phi_{y}\rangle\right\}_{y}$ is an orthonormal set. Consider that
$\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]$
$=\alpha\sum_{y}\omega_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle+\left(1-\alpha\right)\sum_{y}\omega_{y}^{\frac{\alpha}{\alpha-1}}\langle\phi_{y}|\sigma|\phi_{y}\rangle$ (105)
$=\sum_{y}\alpha\omega_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle+\left(1-\alpha\right)\omega_{y}^{\frac{\alpha}{\alpha-1}}\langle\phi_{y}|\sigma|\phi_{y}\rangle$ (106)
$\geq\sum_{y}\left[\omega_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle\right]^{\alpha}\left[\omega_{y}^{\frac{\alpha}{\alpha-1}}\langle\phi_{y}|\sigma|\phi_{y}\rangle\right]^{1-\alpha}$ (107)
$=\sum_{y}\omega_{y}^{\alpha}\omega_{y}^{-\alpha}\langle\phi_{y}|\rho|\phi_{y}\rangle^{\alpha}\langle\phi_{y}|\sigma|\phi_{y}\rangle^{1-\alpha}$ (108)
$=\sum_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle^{\alpha}\langle\phi_{y}|\sigma|\phi_{y}\rangle^{1-\alpha}.$ (109)
The inequality follows from the inequality of weighted arithmetic and geometric means (i.e., $\alpha b+\left(1-\alpha\right)c\geq b^{\alpha}c^{1-\alpha}$ for all $b,c\geq 0$ and $\alpha\in\left(0,1\right)$), applied for all $y$. This inequality is saturated when the following condition holds
$\omega_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle=\omega_{y}^{\frac{\alpha}{\alpha-1}}\langle\phi_{y}|\sigma|\phi_{y}\rangle.$ (110)
That is, it is saturated when the two terms being averaged are equal. Rearranging this, we find that saturation holds if
$\omega_{y}=\left(\frac{\langle\phi_{y}|\sigma|\phi_{y}\rangle}{\langle\phi_{y}|\rho|\phi_{y}\rangle}\right)^{1-\alpha}.$ (111)
Thus, it follows that for every projective measurement $\left(|\phi_{y}\rangle\!\langle\phi_{y}|\right)_{y}$, there exists $\omega>0$ such that
$\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]=\sum_{y}\langle\phi_{y}|\rho|\phi_{y}\rangle^{\alpha}\langle\phi_{y}|\sigma|\phi_{y}\rangle^{1-\alpha}.$ (112)
As such we conclude the desired equality for all $\alpha\in\left(0,1\right)$:
$\inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right\}=Q_{\alpha}^{P}(\rho\|\sigma).$ (113)
Then applying (34) leads to the claim for $\alpha\in\left(0,1\right)$.
The case $\alpha>1$ follows from a very similar proof, but instead makes use of the following inequality: $\alpha b+\left(1-\alpha\right)c\leq b^{\alpha}c^{1-\alpha}$, which holds for all $b\geq 0$, $c>0$, and $\alpha>1$. This inequality is a consequence of Bernoulli’s inequality, which states that $1+rx\leq\left(1+x\right)^{r}$ holds for all $r\geq 1$ and $x\geq-1$. Indeed, consider that
$\alpha b+\left(1-\alpha\right)c\leq b^{\alpha}c^{1-\alpha}\quad$ $\Longleftrightarrow\quad\alpha\left(\frac{b}{c}\right)+\left(1-\alpha\right)\leq\left(\frac{b}{c}\right)^{\alpha}$ (114)
$\Longleftrightarrow\quad\alpha\left(\frac{b}{c}-1\right)+1\leq\left(\frac{b}{c}-1+1\right)^{\alpha},$ (115)
so that we choose $x=\frac{b}{c}-1\geq-1$ and $r=\alpha\geq 1$ in Bernoulli’s inequality. So then we conclude the following for $\alpha>1$:
$\sup_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right\}=Q_{\alpha}^{P}(\rho\|\sigma),$ (116)
by applying the same reasoning as in (104)–(112), but the inequality in (107) goes in the opposite direction for $\alpha>1$.
Appendix B $\alpha\to 1$ limit of the measured Rényi relative entropy
Proof of Proposition 4. This follows because
$\lim_{\alpha\rightarrow 1^{-}}D_{\alpha}^{M}(\rho\|\sigma)$ $=\sup_{\alpha\in\left(0,1\right)}D_{\alpha}^{M}(\rho\|\sigma)$ (117)
$=\sup_{\alpha\in\left(0,1\right)}\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}D_{\alpha}(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (118)
$=\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\sup_{\alpha\in\left(0,1\right)}D_{\alpha}(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (119)
$=\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}D(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (120)
$=D^{M}(\rho\|\sigma),$ (121)
as noted in [RSB24, Lemma 2], and because
$\lim_{\alpha\rightarrow 1^{+}}D_{\alpha}^{M}(\rho\|\sigma)$ $=\inf_{\alpha>1}D_{\alpha}^{M}(\rho\|\sigma)$ (122)
$=\inf_{\alpha>1}\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}D_{\alpha}(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (123)
$=\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}\inf_{\alpha>1}D_{\alpha}(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (124)
$=\sup_{\mathcal{X},\left(\Lambda_{x}\right)_{x\in\mathcal{X}}}D(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ (125)
$=D_{\alpha}^{M}(\rho\|\sigma).$ (126)
The third equality above is non-trivial and follows from the fact that it suffices to optimize $D_{\alpha}(\left(\operatorname{Tr}[\Lambda_{x}\rho]\right)_{x\in\mathcal{X}}\|\left(\operatorname{Tr}[\Lambda_{x}\sigma]\right)_{x\in\mathcal{X}})$ over POVMs with a finite number of outcomes (a compact and convex set) [BFT17, Theorem 4], that the relative entropy $D_{\alpha}$ is lower semi-continuous [vEH14, Theorem 15], and an application of the Mosonyi–Hiai minimax theorem [MH11, Lemma A.2].